The elevator and its passenger together have a mass of 1200.0 kg. While moving upwards, the elevator slows down from 10 m/s to 0 m/s in 2.5 seconds. How much tension (in N) must be in the cable during that time? (Use g=10N/kg)


1
Expert's answer
2022-03-29T05:33:29-0400

The deceleration of the elevator

"a=\frac{\Delta v}{\Delta t}=\frac{0-10}{2.5}=-4.0\:\rm m/s^2"

The tension in the cable

"T=m(g+a)\\
=1200(9.8-4.0)=7000\:\rm N"